# Breathing Black Hole Shadows in Modified Gravity (MOG)

### v1

@@ IV Massive Vector Field Perturbations: The Delayed Echo
The scalar breathing mode derived in Section III provides a compelling, simultaneous signature for the modified gravitational degrees of freedom in Scalar-Tensor-Vector Gravity (STVG). However, the most distinct feature of this theory is its massive vector field, denoted as $\phi_{\mu}$ [12]. In standard General Relativity, gravitational waves are mediated exclusively by a massless tensor field, propagating exactly at the speed of light ($v=c\equiv 1$). Consequently, any transient geometric distortion to a black hole shadow occurs synchronously with the arrival of the electromagnetic signal. In MOG, the massive nature of the vector field introduces dispersion, fundamentally altering this causality timeline and creating a measurable time delay.
To quantify this delayed “echo,” we must strictly analyze the propagation of the vector wave from a distant astrophysical source at a propagation distance $D$. In the STVG action, the massive vector field obeys a Proca-type equation of motion. In a flat or weakly curved background, this is expressed as:
@@
where $\omega$ is the angular frequency and $k$ is the wave number. The velocity of the propagating energy—the group velocity $v_{g}$—is derived by differentiating this dispersion relation:
@@
Unlike transverse-traceless modes, a massive vector field propagating along the line of sight ($z$-axis) heavily excites longitudinal metric polarizations, specifically the $xz$ and $yz$ components. To determine the effect on the shadow, we evaluate the perturbed Hamiltonian for the null geodesics, $H=H_{0}+\delta H=0$. The longitudinal perturbations introduce cross-terms coupling the line-of-sight momentum ($p_{z}$) with the transverse image-plane momenta ($p_{x},p_{y}$):
@@
To rigorously assess the theoretical observability of this secondary delayed echo, we must mathematically evaluate the amplitude of the longitudinal metric perturbation, $h_{\mu\nu}^{\text{vector}}$. In standard classical field theory, the massive vector field $\phi_{\mu}$ contributes to the spacetime geometry via its stress-energy tensor, defined in the MOG action as:
@@
However, this quadratic suppression is partially mitigated by the unique gravitational coupling framework of STVG. In MOG, the vector field is not a standard-model remnant but a primary fifth-force carrier whose effective charge is governed by the deformation parameter, $Q_{G}=\sqrt{\alpha(1+\alpha)}M$. In the highly relativistic environment of an Extreme Mass Ratio Inspiral (EMRI) operating near the photon sphere, the local kinetic energy density of the vector wave is magnified by extreme spatial and temporal gradients, $(\partial\delta\phi)^{2}\sim\omega_{v}^{2}\delta\phi^{2}$, where the quasinormal mode frequency $\omega_{v}$ is significantly high [29]. Consequently, the local effective source term becomes:
@@
where $v\sim\mathcal{O}(1)$ is the orbital velocity of the infalling mass. The direct proportionality to the $\alpha$ parameter demonstrates that the massive vector field carries a macroscopic fraction of the system’s radiated binding energy. While mathematically remaining a second-order metric perturbation, the $\alpha$-enhanced kinetic energy density in the strong-field regime ensures that the resulting translational wobble, though parametrically smaller than the primary scalar breathing mode, retains sufficient amplitude to serve as a physically viable, isolated phenomenological signature.
Having established the physical amplitude of this secondary vector echo, we must also briefly address the theoretical concern of gauge invariance before translating the perturbed Hamiltonian cross-terms into observable celestial coordinates. In standard General Relativity, longitudinal metric perturbations such as $h_{xz}$ and $h_{yz}$ can frequently be eliminated via a suitable coordinate transformation ($x^{\mu}\rightarrow x^{\mu}+\xi^{\mu}$). In such cases, any resulting spatial translations of test particles are non-physical gauge artifacts. However, this is fundamentally not the case in Scalar-Tensor-Vector Gravity. The massive vector field $\phi_{\mu}$ is governed by a Proca-type equation; the presence of the mass term $\mu_{v}$ explicitly breaks the standard local gauge invariance associated with massless fields. This massive force carrier intrinsically possesses longitudinal polarization states. Therefore, the secondary metric modes $h_{xz}$ and $h_{yz}$ sourced by the vector wave represent true, physical degrees of freedom that cannot be globally gauged away by a spatial diffeomorphism. Consequently, the resulting deviations in the transverse photon momenta strictly map to gauge-invariant astrometric shifts. When these trajectories are projected onto the distant observer’s screen, the asymmetric wobble of the shadow’s center represents a genuine, measurable physical displacement rather than a mere coordinate illusion.
@@
To explicitly evaluate these integrals, we apply the standard thin-lens approximation. The photon predominantly accumulates deflection within the highly localized strong-field regime of the effective potential barrier. We define this interaction region by the characteristic diameter of the photon sphere, $\Delta z\approx 2r_{p}$. Furthermore, because the astrophysical wavelength of the massive vector mode is much larger than the black hole scale ($\lambda_{v}\gg r_{p}$), the phase of the wave is approximately constant across the integration bounds. The integral analytically reduces to the product of the interaction length and the time-delayed metric strain.
@@
This completes the phenomenological timeline of a STVG gravitational wave event. An observer will first detect a violent combination of area preserving tensor stretching and area fluctuating scalar breathing. Following a period of static relaxation, a secondary longitudinal wobble defined by Eq. (37) and Eq. (38) will suddenly displace the shadow. Because the magnitude of this spatial translation is directly proportional to the photon sphere radius $r_{p}$, and the temporal separation $\Delta t$ is dictated by the massive dispersion relation, this delayed echo provides a robust, fully parameterized observational template to test the validity of Scalar-Tensor-Vector Gravity.
@@
To properly contextualize these dynamic geometric shifts as a realistic observational template, we must quantitatively clarify the expected astrophysical magnitudes of these perturbations. If one considers a standard gravitational wave originating from a distant galaxy, the strain amplitude reaching the target black hole is typically $h\sim 10^{-21}$. Detecting a fractional area fluctuation or spatial wobble of this minute magnitude is technologically inconceivable for current instruments like the Event Horizon Telescope.
However, the physical relevance of our phenomenological model becomes highly pronounced in the context of an Extreme Mass Ratio Inspiral (EMRI) [34]. Consider a smaller compact object of mass $m$ orbiting near the photon sphere of a supermassive host black hole of mass $M$, such that $m\ll M$. To calculate the dynamic response of the shadow, we must evaluate the local strain before it decays over cosmic distances. Using the standard quadrupole approximation in geometric units ($G=c=1$), the local strain amplitude generated by the smaller mass is inversely proportional to the local propagation distance $D$ and directly proportional to its orbital kinetic energy:
@@ V Conclusion
In this paper, we have rigorously derived the dynamic interaction between gravitational waves and the shadow of a non-rotating black hole within the framework of Scalar-Tensor-Vector Gravity (STVG), also known as Modified Gravity (MOG) [12]. Our primary objective was to mathematically demonstrate how the unique fields introduced by MOG—specifically a dynamic scalar field and a massive vector field—produce observable, geometric deviations from the predictions of standard General Relativity.
@@
The core of our theoretical work involved perturbing this static background with incoming gravitational waves. By injecting a time-dependent metric perturbation into the Hamiltonian, we proved that the MOG scalar field generates a volumetric breathing mode. Our derivation yielded the exact time-dependent radius of the shadow, $R_{\text{sh}}(t)\approx\bar{R}_{\text{sh}}\left(1+\frac{1}{2}h_{b}(t)\right)$. This equation proves that the massless scalar wave directly modulates the critical impact parameter of the trapped photons. Physically, this means the entire black hole shadow will rhythmically expand and contract. The resulting area fluctuation, $\delta A(t)\approx\pi\bar{R}_{\text{sh}}^{2}h_{b}(t)$, is a smoking-gun signature of scalar-tensor gravity, as standard General Relativity strictly preserves the shadow’s area during a tensor wave perturbation.
Furthermore, we demonstrated that the massive nature of the MOG vector field, $\phi_{\mu}$, introduces a fundamental time delay due to quantum vacuum dispersion. Because the group velocity of this massive field is less than the speed of light ($v_{g}<1$), the vector wave arrives at the observer after a predictable delay, $\Delta t$. When this massive wave interacts with the spacetime geometry, it sources secondary, longitudinal metric perturbations. We showed mathematically that these longitudinal cross-terms in the perturbed Hamiltonian shift the effective angular momentum center of the photons. This results in an asymmetric geometric translation, meaning the shadow experiences a sudden, observable wobble in the celestial $X$ and $Y$ coordinates.
To properly contextualize these dynamic geometric shifts as a realistic observational template, we must quantitatively clarify the expected astrophysical magnitudes of these perturbations. If one considers a standard gravitational wave originating from a distant galaxy, the strain amplitude reaching the target black hole is typically $h\sim 10^{-21}$. Detecting a fractional area fluctuation or spatial wobble of this minute magnitude is technologically inconceivable for current instruments. However, the physical relevance of our phenomenological model becomes highly pronounced in the context of an Extreme Mass Ratio Inspiral (EMRI) [34]. Consider a smaller compact object of mass $m$ orbiting near the photon sphere of a supermassive host black hole of mass $M$, such that $m\ll M$. Using the standard quadrupole approximation, the local strain amplitude generated by the smaller mass is inversely proportional to the local propagation distance $D$ and directly proportional to its orbital kinetic energy:
@@
In summary, we have established a complete mathematical framework showing that a MOG gravitational wave event will produce a sequential geometric signature on a black hole shadow: an initial area-fluctuating breathing mode, a brief relaxation, and a delayed translational wobble. While resolving these high-frequency dynamic shifts currently exceeds the temporal resolution of the Event Horizon Telescope [4], our EMRI derivation places the detection of MOG-specific dynamic shadows firmly within the plausible target parameters of proposed space-based interferometry missions and the next-generation Event Horizon Telescope (ngEHT) [35]. These dynamic shadows offer a uniquely powerful laboratory to probe the strong-field regime and fundamentally test the existence of massive force carriers in gravity.

### v2

@@ IV Massive Vector Field Perturbations: The Delayed Echo
The scalar breathing mode derived in Section III provides a compelling, simultaneous signature for the modified gravitational degrees of freedom in STVG. However, the most distinct feature of this theory is its massive vector field, denoted as $\phi_{\mu}$ [12]. In standard General Relativity, gravitational waves are mediated exclusively by a massless tensor field, propagating exactly at the speed of light ($v=c\equiv 1$). Consequently, any transient geometric distortion to a black hole shadow occurs synchronously with the arrival of the electromagnetic signal. In MOG, the massive nature of the vector field introduces dispersion, fundamentally altering this causality timeline and creating a measurable time delay.
To quantify this delayed echo, we must analyze the propagation of the vector wave from a distant astrophysical source over a distance $D$. In the STVG action, the massive vector field obeys a Proca-type equation of motion. In a flat or weakly curved background, this is expressed as,
@@
where $\omega$ is the angular frequency and $k$ is the wave number. The group velocity $v_{g}$—is derived by differentiating this dispersion relation:
@@
Unlike transverse-traceless modes, a massive vector field propagating along the line of sight, $z$-axis, heavily excites longitudinal metric polarizations, specifically the $xz$ and $yz$ components. To determine the effect on the shadow, we evaluate the perturbed Hamiltonian for the null geodesics, $H=H_{0}+\delta H=0$. The longitudinal perturbations introduce cross-terms coupling the line-of-sight momentum ($p_{z}$) with the transverse image-plane momenta ($p_{x},p_{y}$):
@@
To assess the theoretical observability of this secondary delayed echo, we must evaluate the amplitude of the longitudinal metric perturbation, $h_{\mu\nu}^{\text{vector}}$. In standard classical field theory, the massive vector field $\phi_{\mu}$ contributes to the spacetime geometry via its stress-energy tensor, defined in the MOG action as:
@@
However, this quadratic suppression is partially mitigated by STVG’s unique gravitational coupling framework. In MOG, the vector field is not a standard-model remnant but a primary fifth-force carrier whose effective charge is governed by the deformation parameter, $Q_{G}=\sqrt{\alpha(1+\alpha)}M$. In the highly relativistic environment of an Extreme Mass Ratio Inspiral (EMRI) operating near the photon sphere, the local kinetic energy density of the vector wave is magnified by extreme spatial and temporal gradients, $(\partial\delta\phi)^{2}\sim\omega_{v}^{2}\delta\phi^{2}$, where the quasinormal mode frequency $\omega_{v}$ is significantly high [29]. Consequently, the local effective source term becomes:
@@
where $v\sim\mathcal{O}(1)$ is the orbital velocity of the infalling mass. The direct proportionality to the $\alpha$ parameter demonstrates that the massive vector field carries a macroscopic fraction of the system’s radiated binding energy. While mathematically remaining a second-order metric perturbation, the $\alpha$ enhanced kinetic energy density in the strong-field regime ensures that the resulting translational wobble, though parametrically smaller than the primary scalar breathing mode, retains sufficient amplitude to serve as a physically viable, isolated phenomenological signature.
Having established the physical amplitude of this secondary vector echo, we must also briefly address the theoretical concern of gauge invariance before translating the perturbed Hamiltonian cross-terms into observable celestial coordinates. In standard General Relativity, longitudinal metric perturbations such as $h_{xz}$ and $h_{yz}$ can frequently be eliminated via a suitable coordinate transformation ($x^{\mu}\rightarrow x^{\mu}+\xi^{\mu}$). In such cases, any resulting spatial translations of test particles are non-physical gauge artifacts. However, this is fundamentally not the case in Scalar-Tensor-Vector Gravity. The massive vector field $\phi_{\mu}$ is governed by a Proca-type equation; the presence of the mass term $\mu_{v}$ explicitly breaks the standard local gauge invariance associated with massless fields. This massive force carrier intrinsically possesses longitudinal polarization states. Therefore, the secondary metric modes $h_{xz}$ and $h_{yz}$ sourced by the vector wave represent true, physical degrees of freedom that cannot be globally gauged away by a spatial diffeomorphism. Consequently, the resulting deviations in the transverse photon momenta map strictly onto gauge-invariant astrometric shifts. When these trajectories are projected onto the distant observer’s screen, the asymmetric wobble of the shadow’s center represents a genuine, measurable physical displacement rather than a mere coordinate illusion.
@@
To explicitly evaluate these integrals, we apply the standard thin-lens approximation. The photon predominantly accumulates deflection within the highly localized strong-field regime of the effective potential barrier. We define this interaction region by the characteristic diameter of the photon sphere, $\Delta z\approx 2r_{p}$. Furthermore, because the astrophysical wavelength of the massive vector mode is much larger than the black hole scale ($\lambda_{v}\gg r_{p}$), the phase of the wave is approximately constant across the integration bounds. The integral can be analytically reduced to the product of the interaction length and the time-delayed metric strain.
@@
This completes the phenomenological timeline of an STVG gravitational wave event. An observer will first detect a violent combination of area-preserving tensor stretching and area-fluctuating scalar breathing. Following a period of static relaxation, a secondary longitudinal wobble defined by Eq. (37) and Eq. (38) will suddenly displace the shadow. Because the magnitude of this spatial translation is directly proportional to the photon sphere radius $r_{p}$, and the temporal separation $\Delta t$ is dictated by the massive dispersion relation, this delayed echo provides a robust, fully parameterized observational template to test the validity of Scalar-Tensor-Vector Gravity.
@@
To properly contextualize these dynamic geometric shifts within a realistic observational template, we must quantify the expected astrophysical magnitudes of these perturbations. If one considers a standard gravitational wave originating from a distant galaxy, the strain amplitude reaching the target black hole is typically $h\sim 10^{-21}$. Detecting a spatial wobble of this minute magnitude is technologically infeasible with current instruments such as the Event Horizon Telescope.
However, the physical relevance of our phenomenological model becomes highly pronounced in the context of an Extreme Mass Ratio Inspiral (EMRI) [34]. Consider a smaller compact object of mass $m$ orbiting near the photon sphere of a supermassive host black hole of mass $M$, such that $m\ll M$. To calculate the shadow’s dynamic response, we must evaluate the local strain before it decays over cosmic distances. Using the standard quadrupole approximation in geometric units ($G=c=1$), the local strain amplitude generated by the smaller mass is inversely proportional to the local propagation distance $D$ and directly proportional to its orbital kinetic energy:
@@ V Conclusion
In this paper, we derive the dynamical interaction between gravitational waves and the shadow of a non-rotating black hole within the framework of STVG [12]. Our primary objective was to mathematically demonstrate that the unique fields introduced by MOG, specifically, a dynamic scalar field and a massive vector field, produce observable geometric deviations from the predictions of standard General Relativity.
@@
The core of our theoretical work was to perturb this static background with incoming gravitational waves. By injecting a time-dependent metric perturbation into the Hamiltonian, we proved that the MOG scalar field generates a volumetric breathing mode. Our derivation yielded the exact time-dependent radius of the shadow, $R_{\text{sh}}(t)\approx\bar{R}_{\text{sh}}\left(1+\frac{1}{2}h_{b}(t)\right)$. This equation proves that the massless scalar wave directly modulates the critical impact parameter of the trapped photons. Physically, this means that the entire black hole shadow will expand and contract in a rhythmic manner. The resulting area fluctuation, $\delta A(t)\approx\pi\bar{R}_{\text{sh}}^{2}h_{b}(t)$, is a smoking-gun signature of scalar-tensor gravity, as standard General Relativity strictly preserves the shadow’s area during a tensor wave perturbation.
Furthermore, we demonstrated that the massive nature of the MOG vector field, $\phi_{\mu}$, introduces a fundamental time delay due to quantum vacuum dispersion. Because the group velocity of this massive field is less than the speed of light ($v_{g}<1$), the vector wave arrives at the observer after a predictable delay, $\Delta t$. When this massive wave interacts with the spacetime geometry, it sources secondary, longitudinal metric perturbations. We showed that these longitudinal cross-terms in the perturbed Hamiltonian shift the effective angular-momentum center of the photons. This results in an asymmetric geometric translation, meaning the shadow experiences a sudden, observable wobble in the celestial $X$ and $Y$ coordinates.
To properly contextualize these dynamic geometric shifts within a realistic observational template, we must quantify the expected astrophysical magnitudes of these perturbations. If one considers a standard gravitational wave originating from a distant galaxy, the strain amplitude reaching the target black hole is typically $h\sim 10^{-21}$. Detecting a spatial wobble of this minute magnitude is technologically infeasible with current instruments. However, the physical relevance of our phenomenological model becomes highly pronounced in the context of an EMRI [34]. Consider a smaller compact object of mass $m$ orbiting near the photon sphere of a supermassive host black hole of mass $M$, such that $m\ll M$. Using the standard quadrupole approximation, the local strain amplitude generated by the smaller mass is inversely proportional to the local propagation distance $D$ and directly proportional to its orbital kinetic energy:
@@
In summary, we have established a complete mathematical framework showing that a MOG gravitational wave event will produce a sequential geometric signature on a black hole shadow: an initial area-fluctuating breathing mode, a brief relaxation, and a delayed translational wobble. While resolving these high-frequency dynamic shifts currently exceeds the temporal resolution of the Event Horizon Telescope [4], our EMRI derivation places the detection of MOG-specific dynamic shadows firmly within the plausible target parameters of proposed space-based interferometry missions and the next-generation Event Horizon Telescope (ngEHT) [35]. These dynamic shadows provide a uniquely powerful laboratory for probing the strong-field regime and for fundamentally testing the existence of massive force carriers in gravity.